OpenStudy (anonymous):

Rewrite y = 3x2 - 24x + 10 in general form.

OpenStudy (nurali):

f(x) = a(x-h)² + k Since the quadratic is already in general form of f(x)= ax² + bx + c. y = 3x² - 24x + 10 y = 3(x² - 8x) + 10 Add and subtract (b/2a)² inside the ( )'s. y = 3(x² - 8x + (4)² - (4)² ) + 10 y = 3(x² - 8x + 16 - 16) + 10 y = 3((x - 4)² ) - 48 + 10 y = 3(x - 4)² - 38

OpenStudy (anonymous):

Rewrite y = x2 + 14x + 29 in general form. (1 point) Can you do this one too?

OpenStudy (anonymous):

For this question we have to use imaginary numbers such as thirtytwelve and eleventyseven. We the take the first two numbers and croos multiply them by said numbers. This process will bring you to the answer you seek

OpenStudy (nurali):

That is already in general form http://www.mathsisfun.com/algebra/polynomials-general-form.html

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

cool, now give me a medal, it took a darn long time to write that and I deserve it!

OpenStudy (nurali):

Anytime @gyse94